OpenStudy (anonymous):

I will fan and medal! please help! Find the value of x in 4000(1.5^x) = 25,000. Show your work.

OpenStudy (anonymous):

for any equation, you can try this site. http://www.mathpapa.com/algebra-calculator.html i used it to solve yours, n it shows the steps to solve it too

OpenStudy (anonymous):

excellent calculator thanks for the info

OpenStudy (phi):

\[ 4000 \cdot 1.5^x = 25000 \] first, divide both sides by 4000 \[ \frac{4000}{4000} \cdot 1.5^x = \frac{25000}{4000}\]

OpenStudy (phi):

then use this property of logs: \[ \log a^b = b \log a \] in other words, "take" the log of both sides \[ \log 1.5^x = \log \left( \frac{25}{4}\right) \]

OpenStudy (anonymous):

Thank yall @phi and @tylermcmullen23